# Contraherent cosheaves of contramodules on Noetherian formal schemes

### 29 Mar 2026 v1

@@ 2.10. Cotorsion contramodules
Recall that a contramodule $R$-module $\mathfrak{F}$ is flat if and only if it is flat as an $R$-module (or as an $\mathfrak{R}$-module), by Lemma 2.5.10. Furthermore, the inclusion of abelian categories $R{\operatorname{\mathsf{--Contra}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ induces isomorphisms of all the Ext groups/modules, by Proposition 2.5.12. In view of these observations, the equivalence (1) $\Longleftrightarrow$ (2) is provided by [17, Corollary D.2.8(c)]. The equivalence (1) $\Longleftrightarrow$ (3) is a particular case of (1) $\Longleftrightarrow$ (2). ∎
@@ 2.11. The coextension of scalars is quotseparated
This is our version of [32, Theorem 9.5]. The “only if” assertion holds by Lemma 2.6.1(d); the nontrivial implication is the “if”.
@@
The following argument is a special case of [32, proof of Theorem 10.4]. The $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is a contramodule by Lemma 2.6.7(b); we only have to prove that it is a quotseparated contramodule. By Proposition 2.10.2, there exists a short exact sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{C}\longrightarrow\mathfrak{F}\longrightarrow\mathfrak{P}\longrightarrow 0$ with a flat quotseparated contramodule $R$-module $\mathfrak{F}$ and a cotorsion quotseparated contramodule $R$-module $\mathfrak{C}$. By Lemma 2.10.5 applied to the $R$-module $F=S$, the short sequence of $S$-modules $0\longrightarrow\operatorname{Hom}_{R}(S,\mathfrak{C})\longrightarrow\operatorname{Hom}_{R}(S,\mathfrak{F})\longrightarrow\operatorname{Hom}_{R}(S,\mathfrak{P})\longrightarrow 0$ is exact. Now the contramodule $R$-module $\mathfrak{F}$ is separated by Lemma 2.5.6, hence the contramodule $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is separated by Lemma 2.11.4. Thus the contramodule $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is a quotient $S$-module of a separated contramodule $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{F})$, as desired. ∎
@@ 2.12. Colocalization of contraadjustedness and colocality of cotorsion
The left adjoint functor $f_{\sharp}$ is exact. The object $\mathfrak{P}\in R{\operatorname{\mathsf{--Contra}}}$ is adjusted to the right adjoint functor $f^{!}$ in the following sense. By Lemma 2.9.11 applied to the $R$-module $F=S$, the short sequence of contramodule $S$-modules $0\longrightarrow f^{!}\mathfrak{P}\longrightarrow f^{!}\mathfrak{M}\longrightarrow f^{!}\mathfrak{N}\longrightarrow 0$ is exact for any short exact sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$. By [26, Lemma 1.7(b)], for any quotseparated contramodule $R$-module $\mathfrak{P}$ and any quotseparated contramodule $S$-module $\mathfrak{G}$, we have a natural injective map of abelian group
@@
which is an isomorphism when $\mathfrak{P}$ is contraadjusted by [26, Lemma 1.7(e)]. Now if the quotseparated contramodule $S$-module $\mathfrak{G}$ is very flat, then so is the quotseparated contramodule $R$-module $f_{\sharp}\mathfrak{G}$, by Lemma 2.9.9(b). Using the criterion of Lemma 2.9.3(2), one concludes that the functor $f^{!}$ takes contraadjusted quotseparated contramodule $R$-modules to contraadjusted quotseparated contramodule $S$-modules. ∎
@@
Part (a) is obvious. Part (b) is a very basic example of an assertion from the class of results known as “contramodule Nakayama lemmas”. See [20, Lemma 4.2 and Remark 4.3], which are applicable by [20, Theorems 3.3 and 4.1]. A simple argument by induction on the number of generators of $I$, using the passage from $\mathfrak{P}$ to $\mathfrak{P}/s\mathfrak{P}$ (where $s\in I$ is one of the generators), is also possible. ∎
@@
as one can see from Lemma 2.6.3(a). Now (7) is the homological Čech complex [17, formula (1.3) from Section 1.2] for the contraadjusted $\overline{R}$-module $\overline{K}$ with respect to the open covering $\bar{f}_{\alpha}\colon\overline{R}\longrightarrow\overline{S}_{\alpha}$ of the (abstract, nontopological) commutative ring $\overline{R}$. By [17, Lemma 1.2.6(b)], the complex (7) is exact, and by Lemma 2.12.6(a) above it follows that the complex (6) is exact as well. ∎
@@ 2.13. Colocality of exactness
As mentioned in Sections 2.9 and 2.10, the full subcategories $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ and $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ are closed under extensions, cokernels of monomorphisms, and infinite products in $R{\operatorname{\mathsf{--Contra}}}$. (The full subcategory $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ is even closed under quotients.) Thus these full subcategories inherit exact category structures from the abelian exact structure on $R{\operatorname{\mathsf{--Contra}}}$. We will always view $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ and $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ as exact categories with the exact structures inherited from (the abelian exact structure on) $R{\operatorname{\mathsf{--Contra}}}$.
@@
For a generalization, see [25, Lemma 4.21], which is applicable in view of [25, Section 2.4(6)]. ∎
@@
For a generalization, see [25, Lemma 4.13], which is applicable in view of [25, Example 3.8(2)]. ∎
@@
Part (b): the “only if” assertion follows from part (a). To prove the “if”, notice that the map $\mathfrak{F}\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{F})$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ by Lemmas 2.12.7 and 2.5.9(a). If the map $f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{G})$ is an admissible monomorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$, then the map $f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{G})$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ by Lemma 2.6.9(d) for every $\alpha$. Hence the composition $\mathfrak{F}\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{G})$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$. The latter composition is equal to the composition $\mathfrak{F}\longrightarrow\mathfrak{G}\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{G})$, and it follows that $\mathfrak{F}\longrightarrow\mathfrak{G}$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ by the dual version of [4, Proposition 7.6].
@@
Part (a) is a particular case of Lemma 2.13.6(a) (the result of Lemma 2.9.6 is presumed). The proof of part (b) is similar to that of Lemma 2.13.6(b). One needs to use Lemmas 2.9.6 and 2.9.9(b). Part (c) is similar to Lemma 2.13.6(c). The proof of part (d) is similar to that of Lemma 2.13.6(d). One needs to use the criterion of Lemma 2.13.3(2). ∎
@@
This lemma is dual-analogous to Lemmas 2.13.6(a,b,d) and 2.13.7(a,b,d). The “only if” assertions in all parts (a–c) are obvious. Let us prove the “if”.

### 17 Apr 2026 v2

@@ 2.13. Cotorsion contramodules
Recall that a contramodule $R$-module $\mathfrak{F}$ is flat if and only if it is flat as an $R$-module (or as an $\mathfrak{R}$-module), by Lemma 2.5.10. Furthermore, the inclusion of abelian categories $R{\operatorname{\mathsf{--Contra}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ induces isomorphisms of all the Ext groups/modules, by Proposition 2.5.12. In view of these observations, the equivalence (1) $\Longleftrightarrow$ (2) is provided by [18, Corollary D.2.8(c)]. The equivalence (1) $\Longleftrightarrow$ (3) is a particular case of (1) $\Longleftrightarrow$ (2). ∎
2.14. Locality of flatness and very flatness of morphisms
This section is a continuation of Sections 2.102.11.
The following lemma is to be compared with Lemma 2.1.8.
Lemma 2.14.1.
Let $f\colon R\longrightarrow S$ be a taut continuous map of adic topological rings. Then the following two conditions are equivalent:
  1. (1)
    there exists an ideal of definition $I\subset R$ such that, denoting by $J_{n}$ the closure of the ideal $Sf(I^{n})\subset S$, the $R/I^{n}$-module $S/J_{n}$ is flat for every $n\geq 1$;
  2. (2)
    for every open ideal $I\subset R$, denoting by $J$ the closure of the ideal $Sf(I)\subset S$, the $R/I$-module $S/J$ is flat.
Proof.
(2) $\Longrightarrow$ (1) Obvious.
(1) $\Longrightarrow$ (2) Let $I^{\prime}\subset I$ be an open ideal such that, denoting by $J^{\prime}$ the closure of the ideal $Sf(I^{\prime})\subset S$, the $R/I^{\prime}$-module $S/J^{\prime}$ is flat. We need to show that the $R/I$-module $S/J$ is flat whenever $I^{\prime}\subset I$. Indeed, $J^{\prime}$ is an open ideal in $S$ by Lemma 2.10.1(5). Since $J^{\prime}\subset J$, it follows that $J=Sf(I)+J^{\prime}$. Thus $S/J\simeq R/I\otimes_{R/I^{\prime}}S/J^{\prime}$, and the assertion follows. ∎
We will say that a taut continuous homomorphism of adic topological rings $f\colon R\longrightarrow S$ is taut-flat if it satisfies the equivalent conditions of Lemma 2.14.1. Notice that a tight continuous map of adic topological rings is taut-flat if and only if it is flat. Using the criterion of Lemma 2.14.1(2), one can easily check that the composition of any two taut-flat continuous ring maps of adic topological rings is taut-flat.
The next lemma is to be compared with Lemma 2.12.8.
Lemma 2.14.2.
Let $f\colon R\longrightarrow S$ be a taut continuous map of adic topological rings. Then the following three conditions are equivalent:
  1. (1)
    for every open ideal $I\subset R$, denoting by $J$ the closure of the ideal $Sf(I)\subset S$, the induced ring map $\bar{f}\colon R/I\longrightarrow S/J$ is very flat;
  2. (2)
    there exists an ideal of definition $I\subset R$ such that, denoting by $J_{n}$ the closure of the ideal $Sf(I^{n})\subset S$, the induced ring map $\bar{f}\colon R/I^{n}\longrightarrow S/J_{n}$ is very flat for every $n\geq 1$;
  3. (3)
    $f$ is taut-flat and there exists an ideal of definition $I\subset R$ such that, denoting by $J$ the closure of the ideal $Sf(I)\subset S$, the induced ring map $\bar{f}\colon R/I\longrightarrow S/J$ is very flat.
Proof.
(1) $\Longrightarrow$ (2) $\Longrightarrow$ (3) Obvious.
(2) $\Longrightarrow$ (1) Let $I^{\prime}\subset I$ be an open ideal such that, denoting by $J^{\prime}$ the closure of the ideal $Sf(I^{\prime})\subset S$, the ring map $R/I^{\prime}\longrightarrow S/J^{\prime}$ is very flat. We need to show that the ring map $R/I\longrightarrow S/J$ is very flat whenever $I^{\prime}\subset I$. Following the proof of Lemma 2.14.1, we have $S/J\simeq R/I\otimes_{R/I^{\prime}}S/J^{\prime}$. It remains to point out that the ring maps $S/J^{\prime}\longrightarrow S/J$ is surjective (so every localization of $S/J$ at an element arises from a localization of $S/J^{\prime}$ at an element), and refer to [18, Lemma 1.2.2(b)].
(3) $\Longrightarrow$ (2) We have $S/J\simeq R/I\otimes_{R/I^{n}}S/J_{n}$, so [18, Lemma 1.7.10(b)] is applicable. ∎
We will say that a taut continuous homomorphism of adic topological rings $f\colon R\longrightarrow S$ is taut-very flat if it satisfies the equivalent conditions of Lemma 2.14.2. Notice that a tight continuous map of adic topological rings is taut-very flat if and only if it is very flat.
Using the criterion of Lemma 2.14.2(1), one can easily check that the composition of any two taut-very flat continuous ring maps of adic topological rings is taut-very flat. The fact that the composition of any two very flat homomorphisms of commutative rings is very flat (see Section 2.12) needs to be used here.
The following lemma is a generalization of Lemmas 2.3.1(b) and 2.12.9.
Lemma 2.14.3.
Let $f\colon R\longrightarrow S$ be a continuous map of adic topological rings. In this context:
(a) the map $f$ is taut-flat if and only if the map $\Lambda(f)$ is flat;
(b) the map $f$ is taut-very flat if and only if the map $\Lambda(f)$ is very flat.
Proof.
First of all, the map $f$ is taut if and only if the map $\Lambda(f)$ is tight, by Corollary 2.10.4 and Proposition 2.10.7. Assume that this is the case. Put $\mathfrak{R}=\Lambda(R)$ and $\mathfrak{S}=\Lambda(S)$, and denote by $\mathfrak{f}=\Lambda(f)$ the induced ring map $\mathfrak{R}\longrightarrow\mathfrak{S}$. Let $I\subset R$ be an open ideal and $\mathfrak{I}\subset\mathfrak{R}$ be the related open ideal, as per the discussion in Section 2.3. Let $J\subset S$ be the closure of the ideal $Sf(I)\subset S$. Then $\mathfrak{J}=\mathfrak{S}\lambda_{S}(J)$ is the open ideal in $\mathfrak{S}$ corresponding to $J$. Clearly, $\mathfrak{J}$ is the closure of the ideal $\mathfrak{S}\mathfrak{f}(\mathfrak{I})\subset\mathfrak{S}$. Since the map $\mathfrak{f}$ is tight, it follows that $\mathfrak{J}=\mathfrak{S}\mathfrak{f}(\mathfrak{I})$. Now we have $R/I\simeq\mathfrak{R}/\mathfrak{I}$ and $S/J\simeq\mathfrak{S}/\mathfrak{J}$. Both the assertions of parts (a) and (b) follow. ∎
Lemma 2.14.4.
Let $f\colon R\longrightarrow S$ be a continuous homomorphism of adic topological rings. Let $g_{\alpha}\colon S\longrightarrow T_{\alpha}$ be a formal open covering. In this context:
(a) the map $f$ is taut-flat if and only if the composition $g_{\alpha}f\colon R\longrightarrow T_{\alpha}$ is taut-flat for every index $\alpha$;
(b) the map $f$ is taut-very flat if and only if the composition $g_{\alpha}f\colon R\longrightarrow T_{\alpha}$ is taut-very flat for every index $\alpha$.
Proof.
The “only if” assertions hold, since the maps $g_{\alpha}$ are taut-very flat, the compositions of taut-flat morphisms are taut-flat, and the compositions of taut-very flat morphisms are taut-very flat. To prove the “if”, we use the notation from the proof of Proposition 2.11.2. According to that proof, we have $J_{n}=Sf(I^{n})+J_{n+1}\subset S$, hence $J_{n}$ is the closure of the ideal $Sf(I^{n})\subset S$ by the proof of Lemma 2.10.6 (2) $\Rightarrow$ (1). We also know that the ring homomorphisms $\bar{g}_{\alpha}$ form an open covering.
Part (a): we know that $\overline{T}_{\alpha}$ are flat $\overline{R}$-modules, and need to prove that $\overline{S}$ is a flat $\overline{R}$-module. This is a standard fact from commutative algebra; see, e. g., [14, Lemma Tag 01U5] or [18, Lemma 1.8.2].
Part (b): we know that the ring homomorphisms $\bar{g}_{\alpha}\bar{f}$ are very flat, and need to prove that the ring homomorphism $\bar{f}$ is very flat. This result can be found in [18, Lemma 1.8.6]. ∎
Corollary 2.14.5.
Let $f\colon R\longrightarrow S$ be a continuous homomorphism of adic topological rings. Let $g_{\alpha}\colon S\longrightarrow T_{\alpha}$ be a formal open covering. In this setting:
(a) the map $\Lambda(f)$ is flat (as a map of adic topological rings) if and only if the map $\Lambda(g_{\alpha}f)$ is flat for every index $\alpha$;
(a) the map $\Lambda(f)$ is very flat (as a map of adic topological rings) if and only if the map $\Lambda(g_{\alpha}f)$ is very flat for every index $\alpha$.
Proof.
Follows from Lemmas 2.14.3 and 2.14.4. ∎
Lemma 2.14.6.
Let $R$, $S$, and $T$ be adic topological rings, and let $f\colon R\longrightarrow S$ and $g\colon R\longrightarrow T$ be a continuous ring maps. Put $W=T\otimes_{R}S$, and denote by $f^{\prime}\colon T\longrightarrow W$ and $g^{\prime}\colon S\longrightarrow W$ the induced ring maps. Endow the commutative ring $W$ with the tensor product topology, as in Section 2.7. In this setting:
(a) if the ring map $f$ is taut-flat, then the continuous ring map $f^{\prime}$ is also taut-flat;
(b) if the ring map $f$ is taut-very flat and the ring map $g$ is a formal open immersion, then the continuous ring map $f^{\prime}$ is also taut-very flat.
Proof.
Part (a): by Lemma 2.14.3(a), the continuous ring map $\Lambda(f)\colon\Lambda(R)\longrightarrow\Lambda(S)$ is flat. Hence, by Lemma 2.7.3(b), the ring map $\Lambda(T)\longrightarrow\Lambda(T)\otimes_{\Lambda(R)}\Lambda(S)$ is flat, and consequently, taut-flat. By Lemma 2.7.4, we have $\Lambda(W)=\Lambda(\Lambda(T)\otimes_{\Lambda(R)}\Lambda(S))$. Using Lemma 2.14.3(a) again, we conclude that the ring map $\Lambda(f^{\prime})\colon\Lambda(T)\longrightarrow\Lambda(W)$ is flat. By the same lemma, it follows that the ring map $f^{\prime}$ is taut-flat.
The proof of part (b) is similar and based on Lemmas 2.14.3(b) and 2.12.13. ∎
Lemma 2.14.7.
Let $f\colon R\longrightarrow S$ be a continuous homomorphism of adic topological rings, and let $g_{\alpha}\colon R\longrightarrow T_{\alpha}$ be a formal open covering. Put $W_{\alpha}=T_{\alpha}\otimes_{R}S$, and endow the rings $W_{\alpha}$ with the tensor product topologies. Denote by $f^{\prime}_{\alpha}\colon T_{\alpha}\longrightarrow W_{\alpha}$ and $g^{\prime}_{\alpha}\colon S\longrightarrow W_{\alpha}$ the induced ring maps. In this context:
(a) the map $f$ is taut-flat if and only if the map $f^{\prime}_{\alpha}$ is taut-flat for every index $\alpha$;
(b) the map $f$ is taut-very flat if and only if the map $f^{\prime}_{\alpha}$ is taut-very flat for every index $\alpha$.
Proof.
The “only if” assertions hold by Lemma 2.14.6. To prove the “if”, we use the notation from the proof of Proposition 2.11.5. According to that proof, we have $J_{n}=Sf(I^{n})+J_{n+1}\subset S$, hence $J_{n}$ is the closure of the ideal $Sf(I^{n})\subset S$ by the proof of Lemma 2.10.6 (2) $\Rightarrow$ (1).
Part (a): we know that $\overline{W}_{\alpha}$ are flat $\overline{T}_{\alpha}$-modules, and need to prove that $\overline{S}$ is a flat $\overline{R}$-module. This is a standard fact from commutative algebra; see, e. g., [14, Lemma Tag 01U5] or [18, Lemma 1.8.3].
Part (b): we know that the ring homomorphisms $\bar{f}^{\prime}_{\alpha}$ are very flat, and need to prove that the ring homomorphism $\bar{f}$ is very flat. This follows from [18, Lemma 1.2.6(a)]. ∎
Corollary 2.14.8.
Let $f\colon R\longrightarrow S$ be a continuous homomorphism of adic topological rings, and let $g_{\alpha}\colon R\longrightarrow T_{\alpha}$ be a formal open covering. Put $W_{\alpha}=T_{\alpha}\otimes_{R}S$, and endow the rings $W_{\alpha}$ with the tensor product topologies. Denote by $f^{\prime}_{\alpha}\colon T_{\alpha}\longrightarrow W_{\alpha}$ and $g^{\prime}_{\alpha}\colon S\longrightarrow W_{\alpha}$ the induced ring maps. Then the map $\Lambda(f)$ is flat (as a map of adic topological rings) if and only if the map $\Lambda(f^{\prime}_{\alpha})$ is flat for every index $\alpha$.
Proof.
Follows from Lemmas 2.14.3 and 2.14.7. ∎
@@ 2.15. The coextension of scalars is quotseparated
This is our version of [33, Theorem 9.5]. The “only if” assertion holds by Lemma 2.6.1(d); the nontrivial implication is the “if”.
@@
The following argument is a special case of [33, proof of Theorem 10.4]. The $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is a contramodule by Lemma 2.6.7(b); we only have to prove that it is a quotseparated contramodule. By Proposition 2.13.2, there exists a short exact sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{C}\longrightarrow\mathfrak{F}\longrightarrow\mathfrak{P}\longrightarrow 0$ with a flat quotseparated contramodule $R$-module $\mathfrak{F}$ and a cotorsion quotseparated contramodule $R$-module $\mathfrak{C}$. By Lemma 2.13.5 applied to the $R$-module $F=S$, the short sequence of $S$-modules $0\longrightarrow\operatorname{Hom}_{R}(S,\mathfrak{C})\longrightarrow\operatorname{Hom}_{R}(S,\mathfrak{F})\longrightarrow\operatorname{Hom}_{R}(S,\mathfrak{P})\longrightarrow 0$ is exact. Now the contramodule $R$-module $\mathfrak{F}$ is separated by Lemma 2.5.6, hence the contramodule $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is separated by Lemma 2.15.4. Thus the contramodule $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{P})$ is a quotient $S$-module of a separated contramodule $S$-module $\operatorname{Hom}_{R}(S,\mathfrak{F})$, as desired. ∎
@@ 2.16. Colocalization of contraadjustedness and colocality of cotorsion
The left adjoint functor $f_{\sharp}$ is exact. The object $\mathfrak{P}\in R{\operatorname{\mathsf{--Contra}}}$ is adjusted to the right adjoint functor $f^{!}$ in the following sense. By Lemma 2.12.12 applied to the $R$-module $F=S$, the short sequence of contramodule $S$-modules $0\longrightarrow f^{!}\mathfrak{P}\longrightarrow f^{!}\mathfrak{M}\longrightarrow f^{!}\mathfrak{N}\longrightarrow 0$ is exact for any short exact sequence of quotseparated contramodule $R$-modules $0\longrightarrow\mathfrak{P}\longrightarrow\mathfrak{M}\longrightarrow\mathfrak{N}\longrightarrow 0$. By [27, Lemma 1.7(b)], for any quotseparated contramodule $R$-module $\mathfrak{P}$ and any quotseparated contramodule $S$-module $\mathfrak{G}$, we have a natural injective map of abelian group
@@
which is an isomorphism when $\mathfrak{P}$ is contraadjusted by [27, Lemma 1.7(e)]. Now if the quotseparated contramodule $S$-module $\mathfrak{G}$ is very flat, then so is the quotseparated contramodule $R$-module $f_{\sharp}\mathfrak{G}$, by Lemma 2.12.10(b). Using the criterion of Lemma 2.12.3(2), one concludes that the functor $f^{!}$ takes contraadjusted quotseparated contramodule $R$-modules to contraadjusted quotseparated contramodule $S$-modules. ∎
@@
Part (a) is obvious. Part (b) is a very basic example of an assertion from the class of results known as “contramodule Nakayama lemmas”. See [21, Lemma 4.2 and Remark 4.3], which are applicable by [21, Theorems 3.3 and 4.1]. A simple argument by induction on the number of generators of $I$, using the passage from $\mathfrak{P}$ to $\mathfrak{P}/s\mathfrak{P}$ (where $s\in I$ is one of the generators), is also possible. ∎
@@
as one can see from Lemma 2.6.3(a). Now (7) is the homological Čech complex [18, formula (1.3) from Section 1.2] for the contraadjusted $\overline{R}$-module $\overline{K}$ with respect to the open covering $\bar{f}_{\alpha}\colon\overline{R}\longrightarrow\overline{S}_{\alpha}$ of the (abstract, nontopological) commutative ring $\overline{R}$. By [18, Lemma 1.2.6(b)], the complex (7) is exact, and by Lemma 2.16.6(a) above it follows that the complex (6) is exact as well. ∎
@@ 2.17. Colocality of exactness
As mentioned in Sections 2.12 and 2.13, the full subcategories $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ and $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ are closed under extensions, cokernels of monomorphisms, and infinite products in $R{\operatorname{\mathsf{--Contra}}}$. (The full subcategory $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ is even closed under quotients.) Thus these full subcategories inherit exact category structures from the abelian exact structure on $R{\operatorname{\mathsf{--Contra}}}$. We will always view $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cta}}$ and $R{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$ as exact categories with the exact structures inherited from (the abelian exact structure on) $R{\operatorname{\mathsf{--Contra}}}$.
@@
For a generalization, see [26, Lemma 4.21], which is applicable in view of [26, Section 2.4(6)]. ∎
@@
For a generalization, see [26, Lemma 4.13], which is applicable in view of [26, Example 3.8(2)]. ∎
@@
Part (b): the “only if” assertion follows from part (a). To prove the “if”, notice that the map $\mathfrak{F}\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{F})$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ by Lemmas 2.16.7 and 2.5.9(a). If the map $f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow f_{\alpha}^{\sharp}(\mathfrak{G})$ is an admissible monomorphism in $S_{\alpha}{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$, then the map $f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{G})$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ by Lemma 2.6.9(d) for every $\alpha$. Hence the composition $\mathfrak{F}\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{F})\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{G})$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$. The latter composition is equal to the composition $\mathfrak{F}\longrightarrow\mathfrak{G}\longrightarrow\bigoplus_{\alpha}f_{\alpha}{}_{\sharp}f_{\alpha}^{\sharp}(\mathfrak{G})$, and it follows that $\mathfrak{F}\longrightarrow\mathfrak{G}$ is an admissible monomorphism in $R{\operatorname{\mathsf{--Contra}}}_{\mathsf{fl}}$ by the dual version of [5, Proposition 7.6].
@@
Part (a) is a particular case of Lemma 2.17.6(a) (the result of Lemma 2.12.6 is presumed). The proof of part (b) is similar to that of Lemma 2.17.6(b). One needs to use Lemmas 2.12.6 and 2.12.10(b). Part (c) is similar to Lemma 2.17.6(c). The proof of part (d) is similar to that of Lemma 2.17.6(d). One needs to use the criterion of Lemma 2.17.3(2). ∎
@@
This lemma is dual-analogous to Lemmas 2.17.6(a,b,d) and 2.17.7(a,b,d). The “only if” assertions in all parts (a–c) are obvious. Let us prove the “if”.